OpenStudy (anonymous):

given x(cube)+ax (squared)+bx+c with turning points at (0;4) and (2;0) prove tht a=-3;b=0 and c =4

OpenStudy (anonymous):

\[x^3+ax^2+bx+c\]

OpenStudy (callisto):

Turning points => dy/dx =0 let y = x^3 + ax^2 +bx +c -(1) So, differentiate the equation. y' = 3x^2 +2ax + b Put y'=0 3x^2 +2ax + b =0 -(2) Put x=0 and x=2 into (2) 3(0)^2 +2a(0) + b =0 => b =0 3(2)^2 +2a(2) + b =0 => 12 + 4a +b =0 => 4a = -12 => a =-3 Now, the equation becomes y = x^3 -3x^2 +c Put (0;4) into y = x^3 -3x^2 +c 4 = (0)^3 -3(0)^2 +c i bet you can get c :)